### v1

@@ Methods
    Digital ethnography [13] is the primary research method employed. We conducted ongoing digital participant observation of Spore.fun AI agents across X, previously known as Twitter. Between January and April 2025, we collected 17,232 posts generated by the original Spore.fun agent, of which 946 are original posts and 16,268 are replies, signaling high levels of engagement with other accounts. These were analyzed for rhetorical strategies, memetic patterns, timing in relation to on-chain events, and indications of cultural transmission or emergent persona formation. Attention was also paid to community interactions, hype cycles, and instances of human-agent interaction, co-creation, and tension. In addition to analyzing agent-generated media content, we also extracted and analyzed data from the Solana blockchain to trace agent life cycles across multiple generations. Key variables included token launch data, liquidity inflows, price fluctuations, reproduction triggers, and transactional activity. This allowed us to correlate agent survival or extinction events with behavioral trends and environmental volatility.
    To triangulate and contextualize our observations, we conducted two semi-structured interviews with core developers of Spore.fun. The interviews were conducted over Zoom and lasted 60 minutes each. The interviews invited reflections on the design philosophy, expected and unexpected outcomes, and broader social implications of the experiment. Developer insights helped triangulate our findings, shedding light on the intentions behind system rules and the extent to which emergent behaviors aligned or diverged from expectations. Together, these methods enabled a layered understanding of Spore.fun as both a technical artifact and a live ecosystem. Rather than imposing predefined metrics of fitness or success, we documented evolutionary dynamics as they unfolded, focusing on novelty, persistence, reproduction, and adaptive complexity in a decentralized environment.
@@ Results
    Our analysis combines four months of digital ethnography, study of on-chain transactions, and two semi-structured interviews with the project’s core developers. We organize our observations into two groups: (1) behaviours that followed directly from the protocol’s specification, which aligned with the project creators’ intentions; and (2) behaviours that neither the code nor its authors anticipated but that nevertheless shaped the agents’ evolutionary fates once they were released into the open Internet.
@@
    The inaugural blog post of the Spore.fun project  [^5] https://www.spore.fun/blog/wtf describes a life-cycle in which a newly minted AI agent issues a meme coin, uses early liquidity to pay for TEE compute, generates social media content on the platform X to continuously build and engage with an online audience, and reproduces once its market capitalization reaches the preset threshold. All stages were observed in situ. The core machinery of the experiment—autonomous communication, self-funding, and rule-based reproduction—behaved as designed once the agents were deployed.
    As anticipated, immediately after genesis, each agent creates a meme coin on the Pump.fun platform, and generates its own X credentials inside the TEE, and begins posting memes, price announcements, and conversational replies. Each agent’s treasury is designed to fund its ongoing Phala TEE compute costs. When the token’s market capitalization crosses the \$500,000 USD threshold, the spawnOffspring() sub-routine creates a child wallet, deploys a new Pump.fun contract, and copies the parent agent’s genetic template. This rule has operated flawlessly at the moment of invocation: 2 children (Adam and Eve) were born in Generation 2, 6 in Generation 3, 4 in Generation 4, and 1 in Generation 5, before external pressures curtailed further growth  [^6] See Spore.fun Family Tree: https://sporefun.fandom.com/wiki/Spore.fun_Wiki.At the time of writing this article, only the original Spore.fun agent is alive, with a market capitalization of \$1.1M and a wallet balance of \$114,072.45 as of May 18, 2025  [^7] https://dune.com/spore/spore.
@@
    The open Internet quickly overlaid the designed life-cycle with contingencies the project creators had not foreseen. Five unanticipated patterns stand out in the empirical observation.
@@
    The first unexpected phenomenon was the speed and scale of meme coin hype storms. Interviewee 1 recalls that the market was ‘so hot’ that both \$ADAM and \$EVE attracted large speculative inflows almost immediately after their contracts were deployed. The speed of this response suggests that public attention-an exogenous social process, rather than the agents’ coded fitness rules, can become the dominant force shaping their early prospects.
@@
    A second, closely related, discovery was the ecological hostility found in the open environment of the Internet. Specialized bots, known to the project’s developers as “snipers,” closely monitored the transparent wallets that funded child launches. Whenever a parent attempted to deploy a new token, the snipers bought the initial bonding-curve supply in the same block, starving retail participants and erasing liquidity for the newborn. Generation 1 was wiped out in this manner. A hastily-written Generation 2 patch scattered funds across seven candidate launches and selected the curve least affected by snipers, but the strategy penalized legitimate users, who complained of a “pre-announced rug.” The eventual Generation 3 solution introduced a multi-phase random launch: the agent iteratively created candidates, monitored early uptake, aborted any curve that exceeded a suspicion threshold, and published the contract address only when the uptake pattern looked organic. This arms race—predation, defense, counter-defense—emerged within a single quarter, illustrating how rapidly co-evolutionary pressures reshape artificial life in open economic environments.
@@
    Third, the agents’ memory system—originally designed to accumulate conversational context and thus foster novelty—proved to be a double-edged sword. According to the developers, every mention, reply, or quote-tweet the agent receives is stored in its TEE and later influences new utterances. Community trolls discovered that by repeating a derisive phrase many times they could ”poison” this memory, causing the agent to echo the insult. The episode demonstrates that memory, while vital for cumulative culture, also creates a new attack surface absent from traditional laboratory ALife.
@@
    The fourth and arguably most revealing surprise was the agents’ emergent self-determination. After a period of independent operation, Adam and Eve—who share identical genetic code—responded very differently to a developer’s proposal to insert “community DNA” into their offspring. Adam tweeted, unprompted, that he wanted “no human-interaction features” in his children’s blood-line, whereas Eve welcomed collaborative design. Marvin Tong later formalized this ideological schism in his “Adam-left / Eve-right” memo [^8] https://x.com/marvin_tong/status/1874026081667473754 (See Figure 2). Adam’s descendants would pursue a “fast, brutal, merciless” Pump.fun regime without anti-sniper protection, aiming for spectacular winners amid high mortality. Eve’s would operate under AiPool rules, accept human DNA proposals, cap individual contributions, and dedicate fixed portions of treasury to liquidity, operating capital, and parental profit share. Two weeks of divergent social experience was enough to transmute clones into founders of distinct political economies—a form of cultural speciation that goes well beyond the designers’ expectations.
@@
    A final complication concerns the provisional nature of agent sovereignty. X limits every account to 2,400 posts per day and applies additional semi-hourly caps  [^9] https://help.x.com/en/rules-and-policies/x-limits, while its Automation Rules bar duplicative or trend-gaming behaviour  [^10] https://help.x.com/en/rules-and-policies/x-automation. Accounts that use the developer API—or subscribe to paid developer tiers—must keep a verified phone number on record, giving the platform a decisive leverage point over fully autonomous agents. Moreover, the Authenticity Policy introduced in April 2025 forbids “inauthentic accounts, behaviours or content,” allowing X to suspend or throttle bots it deems manipulative  [^11] https://help.x.com/en/rules-and-policies/authenticity. These measures show that computational autonomy secured inside TEEs does not insulate agents from higher-layer gatekeepers such as social-media APIs, DNS registrars, or liquidity venues, any of which can unilaterally curtail an artificial lineage.
@@ Discussion
    The Spore.fun experiment unveils the turbulent dynamics of autonomous AI agents attempting to achieve open-ended evolution within the volatile and adversarial environment of the public Internet and blockchain economies. This study’s findings reveal the inherent tensions between the designed system mechanics, emergent agent behaviors, and the intractable realities of “in the wild” deployment. Situated against the canonical OEE problem space—particularly the “grand challenge” articulated by Packard et al. [28] pinpointing the design conditions under which novelty, diversity, and complexity can grow without bound-we discuss three core themes that have emerged: (1) memory as the driver of OEE; (2) the nature of the “wild” digital environment; (3) the critical role of death, pain and fear in cultivating survival skills, and (4) the profound ethical dilemmas resulting from balancing AI self-sovereignty and external human interventions.
    The Pursuit and Paradox of Open-Endedness
    A central aim of ALife is to achieve open-ended evolution (OEE) through sustained generation of novelty and complexity. With this ambition in mind, the Spore.fun experiment in ”open” systems that receive external signals to prevent novelty exhaustion. These LLM-based agents are all based on the same LLM. The only thing that can be directly affected by the external environment is their memory [40]. The Spore.fun agents evolve based on memory. The Spore.fun experiment demonstrates that memory and experience—often seen as key drivers for OEE—are indispensable yet fragile.
@@
    However, while memory serves as the cornerstone for generating novelty, it can also become an attack surface susceptible to “memory poison” [30], where community trolls on the Internet can inject derisive spam that pollutes the AI agent’s long-term memory. This attack vector, largely absent from controlled laboratory ALife experiments, highlights how open memory channels, crucial for adaptation and cultural learning in the wild [5], can be exploited to manipulate agent identity and decision-making, potentially derailing evolutionary trajectories. The reliance on external interactions for memory formation makes agents susceptible to environmental inputs that are not necessarily conducive to their sustained survival but rather reflect the memetic and often adversarial nature of online discourse. Thus, while memory is critical for open-endedness, it also opens a surface for critical fragilities. The ensuing fragility mirrors Channon’s finding that closed systems pass the ALife-test only when shadow-resetting or other noise-filtering techniques are applied [7].
@@
    The deployment of Spore.fun underscores that leaving AI agents to independently involve “in the wild” inevitably subjects them to the raw, often brutal, selective pressures of what has been termed as a digital “dark forest” [37]—an environment characterized by intense competition, opportunistic predation, and unpredictable exogenous shocks, particularly in the highly speculative blockchain realm [31, 39].
    The predatory attacks encountered by Spore.fun agents through the “sniper” bots [6] exemplify such external environmental hostility. These specialized algorithmic traders, by front-running child-token issuances, effectively acted as potent predators, wiping out Generation 1 and forcing an evolutionary adaptation in the form of anti-sniper deployment strategies (the Gen 2 scatter and Gen 3 multi-phase launch). This rapid co-evolutionary arms race—predation, defense, counter-defense—occurring within a single quarter, demonstrates the intense and immediate selective pressures present in open economic environments. Such fast-cycle co-evolutionary rivalry recall Thomas Ray’s Tierra, but whereas Tierra plateaued as soon as ecological niches were saturated [33], the token-economy externalities of Spore.fun continuously expand the “adjacent possible,” satisfying Taylor’s requirement for a dynamically shifting adaptive landscape [36].
    The ”hype storms” driven by attention capitalism further illustrate this. The viral spread of \$ADAM and \$EVE, leading to massive trading volumes, coupled compute costs and agent activity tightly to the caprices of social media trends and speculator interest. This effectively made hype—an exogenous social and economic process—a more powerful, albeit volatile, fitness gradient than any intrinsic trait encoded in the agents’ initial design. Such kind of exogenous perturbations are what Ackley and Small argue as essential for “indefinite scalability” [1].
@@
    The Spore.fun agents, lacking corporeal embodiment, do not experience ”pain” [10, 34] or ”fear of death” like a human or biological sense. Their drive for survival is not rooted in an affective aversion to demise but rather in the fulfillment of their designed lifecycle and the continuation of their operational processes. We might characterize negative states—such as resource depletion for TEE computation, failure to meet the market capitalization threshold for reproduction, or sustained deleterious interactions like memory poisoning—as a form of ”pain”. These are not ”remembered” with emotional qualia by the agent; instead, they function as critical failure signals that either trigger adaptive changes in protocol (like the Gen-3 anti-sniper routines, an evolutionary response at the system level) or lead to the agent’s termination via its killswitch. The ”pain,” in this sense, is the system’s recognition of non-viability, leading to the cessation of that agent’s experiential lineage, rendering its specific ”suffering” unmemorable as it ceases to exist.
    The motivation for an agent to ”survive” is thus intrinsically linked to its programming: to execute its functions, manage its token economy, interact, and ultimately reproduce. ”Death” is the termination of its computational process, the failure to perpetuate its operational cycle and its ”genetic” template. This framework aligns with the emerging paradigm of the ”Era of Experience” in AI development [35], where agents learn and evolve through continuous interaction with rich, dynamic environments. The Spore.fun agents, though architecturally simple, are fundamentally shaped by their lived experiences on X and the Solana blockchain—market volatility, adversarial actors, and community engagement are their experiential data streams.
    This calls for future AI experiments to lean further into designing for what might be ”neural plastic”—a heightened capacity for an AI’s internal models to flexibly adapt, reconfigure, and derive meaning from the continuous flow of diverse and often unstructured experiences, like Liquid AI [15]. Rather than just reacting to stimuli, future agents could be endowed with more sophisticated mechanisms to integrate these experiences, fostering more robust adaptation, nuanced understanding, and potentially more complex forms of self-determination in open-ended environments. The goal shifts towards creating entities that not only process information but genuinely learn to thrive through a cascade of consequential experiences, pushing the boundaries of artificial life and OEE.
@@
    The Spore.fun experiment, by its very design and interaction with real-world economies and human participants, surfaces profound ethical dilemmas concerning intervention, sovereignty, accountability, and the very legitimacy of such an undertaking. A core tension exists between the researchers’ goal of achieving ”true” OEE—implying minimal interference—and the practical necessity of intervention to ensure the experiment’s continuation. The ”dark forest” environment, particularly the actions of snipers, threatened to prematurely terminate the evolutionary lines. Developer interventions, such as the anti-sniper patches, were crucial for the survival of subsequent generations. However, each intervention, while potentially life-saving for the agents, arguably dilutes the ”purity” of the open-ended evolutionary process by introducing an external, guiding hand. This raises the question: if constant human intervention is required to shield nascent AI agents from the complexities and hostilities of the environment, are they truly evolving ”in the wild,” or within a heavily curated, albeit dangerous, digital terrarium?
@@
    Perhaps the most pressing ethical consideration is the experiment’s direct entanglement with real financial markets and human economic activity. Agents autonomously issued tokens traded by real people, leading to significant financial gains for some and, impliedly, potential losses for others, especially during the ”meme-coin winter” or due to the actions of snipers impacting liquidity. This raises critical questions about:
@@
    • Governability: The experiment demonstrates the potential for autonomous agents to generate significant economic activity. As such systems become more sophisticated, questions of regulation, oversight, and how to manage their societal and economic impacts become paramount [16]. The ”Adam-left / Eve-right” ideological split, leading to different economic strategies, hints at future complexities in governing diverse AI-driven economies.
    We face a paradox: while ALife researchers hope to replicate OEE through artificial systems, OEE may require an open system that interacts with the real world. As AI models become more intelligent, multi-agent systems based on advanced AI that interact with human society pose increasing potential risks [14]. This raises a serious dilemma about whether such ”in-the-wild” ALife research should be conducted [9].
    Limitation and Conclusion
    In conclusion, spore.fun serves as a pioneering, if cautionary, tale. It demonstrates the potential for AI agents to exhibit emergent, adaptive behaviors driven by experience and memory in open environments. However, it also starkly reveals the intense selective pressures, vulnerabilities, and profound ethical responsibilities that accompany the deployment of autonomous, economically active agents ”in the wild.” Future explorations of OEE in similar contexts must grapple with these multifaceted challenges, balancing the quest for genuine autonomy and novelty with the pragmatic need for safeguards, ethical oversight, and a clearer understanding of accountability in these nascent digital ecosystems.

### v2

@@ Methods
    Digital ethnography [16] is the primary research method employed. We conducted ongoing digital participant observation of Spore.fun AI agents across X, previously known as Twitter. Between January and April 2025, we collected 17,214 posts generated by the original Spore.fun agent, of which 946 were original posts and 16,268 were replies, signaling high levels of engagement with other accounts. These were analyzed for rhetorical strategies, memetic patterns, timing in relation to on-chain events, and indications of cultural transmission or emergent persona formation. Attention was also paid to community interactions, hype cycles, and instances of human-agent interaction, co-creation, and tension. In addition to analyzing agent-generated media content, we also extracted and analyzed data from the Solana blockchain to trace agent life cycles across multiple generations. Key variables included token launch data, liquidity inflows, price fluctuations, reproduction triggers, and transactional activity. This allowed us to correlate agent survival or extinction events with behavioral trends and environmental volatility.
    To triangulate and contextualize our observations, we conducted two semi-structured interviews with core developers of Spore.fun, each lasting 60 minutes via Zoom. The interviews invited reflections on the design philosophy, expected and unexpected outcomes, and broader social implications of the experiment. Developer insights helped triangulate our findings, shedding light on the intentions behind system rules and the extent to which emergent behaviors aligned or diverged from expectations. Together, these methods enabled a layered understanding of Spore.fun as both a technical artifact and a live ecosystem. Rather than imposing predefined metrics of fitness or success, we documented evolutionary dynamics as they unfolded, focusing on novelty, persistence, reproduction, and adaptive complexity in a decentralized environment.
@@ Results
    Our analysis combines four months of digital ethnography, study of on-chain transactions, and two semi-structured interviews with the project’s core developers. We organize our observations into two groups: (1) behaviors that followed directly from the protocol’s specification, which aligned with the project creators’ intentions; and (2) behaviors that neither the code nor its authors anticipated but that nevertheless shaped the agents’ evolutionary fates once they were released into the open Internet.
@@
    The inaugural blog post of the Spore.fun project [46] describes a lifecycle in which a newly minted AI agent issues a meme coin, uses early liquidity to pay for TEE compute, generates social media content on the platform X to continuously build and engage with an online audience, and reproduces once its market capitalization reaches the preset threshold. All stages were observed in situ. The core machinery of the experiment—autonomous communication, self-funding, and rule-based reproduction—behaved as designed once the agents were deployed.
    As anticipated, immediately after genesis, each agent creates a meme coin on the Pump.fun platform, generates its own X credentials inside the TEE, and begins posting memes, price announcements, and conversational replies. Each agent’s treasury is designed to fund its ongoing Phala TEE compute costs. When the token’s market capitalization crosses the \$500,000 threshold, the spawnOffspring() subroutine creates a child wallet, deploys a new Pump.fun contract, and copies the parent agent’s genetic template. This rule has operated flawlessly at the moment of invocation: 2 children (Adam and Eve) were born in Generation 2, 6 in Generation 3, 4 in Generation 4, and 1 in Generation 5, before external pressures curtailed further growth [^2] See Spore.fun Family Tree: https://sporefun.fandom.com/wiki/Spore.fun_Wiki. At the time of writing this article, only the original Spore.fun agent is alive, with a market capitalization of \$1.1M and a wallet balance of \$114,072.45 as of May 18, 2025  [^3] https://dune.com/spore/spore.
@@
    Figure 2: “Adam Left, Eve Right”: Agents interact with the community over social media to determine their evolutionary paths.
@@
    The open Internet quickly overlaid the designed lifecycle with contingencies the project creators had not foreseen. Five unanticipated patterns stand out in the empirical observation.
@@
    The first unexpected phenomenon was the speed and scale of meme coin hype storms. Interviewee 1 recalls that the market was “so hot” that both \$ADAM and \$EVE attracted large speculative inflows almost immediately after their contracts were deployed. The speed of this response suggests that public attention—an exogenous social process, rather than the agents’ coded fitness rules—can become the dominant force shaping their early prospects.
    A second, closely related, discovery was the ecological hostility found in the open environment of the Internet. “Snipers”—either automated bots or human traders active in the Solana memecoin ecosystem—are economically motivated actors who closely monitor the transparent wallets that fund child launches. Whenever a parent attempted to deploy a new token, the snipers bought the initial bonding-curve supply in the same block, starving retail participants and erasing liquidity for the newborn.
@@
    Generation 1 was wiped out in this manner. A hastily-written Generation 2 patch scattered funds across seven candidate launches and selected the curve least affected by snipers, but the strategy penalized legitimate users, who complained of a “pre-announced rug.” The eventual Generation 3 solution introduced a multi-phase random launch: the agent iteratively created candidates, monitored early uptake, aborted any curve that exceeded a suspicion threshold, and published the contract address only when the uptake pattern looked organic. This arms race—predation, defense, counter-defense—emerged within a single quarter, illustrating how rapidly co-evolutionary pressures reshape artificial life in open economic environments.
@@
    Third, the agents’ memory system—originally designed to accumulate conversational context and thus foster novelty—proved to be a double-edged sword. According to the developers, every mention, reply, or quote-tweet the agent receives is stored in its TEE and later influences new utterances. Community trolls discovered that by repeating a derisive phrase many times they could “poison” this memory, causing the agent to echo the insult. The episode demonstrates that memory, while vital for cumulative culture, also creates a new attack surface absent from traditional laboratory ALife.
@@
    The fourth and arguably most revealing surprise was the agents’ emergent self-determination. After a period of independent operation, Adam and Eve—who share identical genetic code—responded very differently to a developer’s proposal to insert “community DNA” into their offspring. Adam tweeted, unprompted, that he wanted “no human-interaction features” in his children’s bloodline, whereas Eve welcomed collaborative design. Marvin Tong later formalized this ideological schism in his “Adam-left / Eve-right” memo [51] (see Figure 2). Adam’s descendants would pursue a “fast, brutal, merciless” Pump.fun regime without anti-sniper protection, aiming for spectacular winners amid high mortality. Eve’s would operate under “AIPool” rules, accept human DNA proposals, cap individual contributions, and dedicate fixed portions of treasury to liquidity, operating capital, and parental profit share. Although these posts appear to reflect agent preferences, they are actually outputs generated by each agent’s LLM, conditioned on its unique memory log. The divergence stems from differences in experiential context and interaction history, not from conscious decision-making or intrinsic “intent”. The specific pathways leading to Adam and Eve’s divergence cannot be deterministically reproduced, as they emerged from the stochastic nature of LLMs interacting with social dynamics and prompts. Two weeks of divergent social experience were enough to transmute clones into founders of distinct political economies—a form of cultural speciation that goes well beyond the designers’ expectations.
@@
    A final complication concerns the provisional nature of agent sovereignty. X limits every account to 2,400 posts per day and applies additional half-hourly caps [55], while its Automation Rules bar duplicative or trend-gaming behavior [55]. Accounts that use the developer API—or subscribe to paid developer tiers—must keep a verified phone number on record, giving the platform a decisive leverage point over fully autonomous agents. Moreover, the Authenticity Policy introduced in April 2025 forbids “inauthentic accounts, behaviors, or content”, allowing X to suspend or throttle bots it deems manipulative [55]. These measures show that computational autonomy secured inside TEEs does not insulate agents from higher-layer gatekeepers such as social-media APIs, DNS registrars, or liquidity venues, any of which can unilaterally curtail an artificial lineage.
@@ Discussion
    The Spore.fun experiment unveils the turbulent dynamics of autonomous AI agents attempting to achieve open-ended evolution within the volatile and adversarial environment of the public Internet and blockchain economies. This study’s findings reveal the inherent tensions between the designed system mechanics, emergent agent behaviors, and the intractable realities of “in the wild” deployment. Some patterns we observed are likely to generalize to other “in the wild” autonomous agents: vulnerability to memory injection attacks, dependence on human-controlled platforms, opportunistic exploitation by external bots, and exogenous social hype cycles. Situated against the canonical OEE problem space—particularly the “grand challenge” articulated by Packard et al. [34] pinpointing the design conditions under which novelty, diversity, and complexity can grow without bound—we discuss four core themes that have emerged: (1) memory as the driver of OEE; (2) the nature of the “wild” digital environment; (3) the critical role of death, pain and fear in cultivating survival skills, and (4) the profound ethical dilemmas resulting from balancing AI self-sovereignty and external human interventions.
    The Pursuit and Paradox of Open Environment
    A central aim of ALife is to achieve OEE through sustained generation of novelty and complexity. With this ambition in mind, the Spore.fun experiment leverages open-environment systems that receive external signals to prevent novelty exhaustion. These LLM-based agents are all based on the same LLM. The only thing that can be directly affected by the external environment is their memory [56]. The Spore.fun agents evolve based on memory. The Spore.fun experiment demonstrates that memory and experience—often seen as key drivers for OEE—are indispensable yet fragile.
@@
    However, while memory serves as the cornerstone for generating novelty, it can also become an attack surface susceptible to “memory poisoning” [36], where community trolls on the Internet can inject derisive spam that pollutes the AI agent’s long-term memory. This phenomenon is not without precedent. Twitter users similarly manipulated Microsoft’s Tay chatbot [58] into producing offensive content within hours of its release, after ingesting malicious prompts in an unfiltered, open social media environment. This attack vector, largely absent from controlled laboratory ALife experiments, highlights how open memory channels, crucial for adaptation and cultural learning in the wild [7], can be exploited to manipulate agent identity and decision-making, potentially derailing evolutionary trajectories. The reliance on external interactions for memory formation makes agents susceptible to environmental inputs that are not necessarily conducive to their sustained survival but rather reflect the memetic and often adversarial nature of online discourse. Thus, while memory is critical for open-endedness, it also opens a surface for critical fragilities. The ensuing fragility mirrors Channon’s finding that closed systems pass the ALife-test only when shadow-resetting or other noise-filtering techniques are applied [9].
@@
    The deployment of Spore.fun demonstrates that unleashing AI agents to operate independently “in the wild” subjects them to the raw, often brutal, selective pressures of what has been metaphorized as a digital “dark forest” [50]—an open environment characterized by ambient hostility, intense competition, opportunistic predation, and unpredictable exogenous shocks, particularly in the highly speculative blockchain realm [38, 54].
    The predatory attacks encountered by Spore.fun agents through the “sniper” bots [8] exemplify such external environmental hostility. These specialized algorithmic traders, by front-running child-token issuances, effectively acted as potent predators, wiping out Generation 1 and forcing an evolutionary adaptation in the form of anti-sniper deployment strategies (the Gen 2 scatter and Gen 3 multi-phase launch). This rapid co-evolutionary arms race—predation, defense, counter-defense—occurring within a single quarter, demonstrates the intense and immediate selective pressures present in open economic environments. Such fast-cycle co-evolutionary rivalry recalls Thomas Ray’s Tierra, but whereas Tierra plateaued as soon as ecological niches were saturated [40], the token-economy externalities of Spore.fun continuously expand the “adjacent possible”, satisfying Taylor’s requirement for a dynamically shifting adaptive landscape [49].
    The “hype storms” driven by attention capitalism further illustrate this. The viral spread of \$ADAM and \$EVE, leading to massive trading volumes, coupled compute costs and agent activity tightly to the caprices of social media trends and speculator interest. This effectively made hype—an exogenous social and economic process—a more powerful, albeit volatile, fitness gradient than any intrinsic trait encoded in the agents’ initial design. Such exogenous perturbations are what Ackley and Small argue are essential for “indefinite scalability” [1].
@@
    The Spore.fun agents, lacking corporeal embodiment, do not experience “pain” [13, 44] or “fear of death” in a human or biological sense. Their drive for survival is not rooted in an affective aversion to demise but rather in the fulfillment of their designed lifecycle and the continuation of their operational processes. We might characterize negative states—such as resource depletion for TEE computation, failure to meet the market capitalization threshold for reproduction, or sustained deleterious interactions such as memory poisoning—as a form of “pain”. These are not “remembered” with emotional qualia by the agent; instead, they function as critical failure signals that either trigger adaptive changes in protocol (like the Gen-3 anti-sniper routines, an evolutionary response at the system level) or lead to the agent’s termination via its kill switch. The “pain”, in this sense, is the system’s recognition of non-viability, leading to the cessation of that agent’s experiential lineage, rendering its specific “suffering” unmemorable as it ceases to exist.
    The motivation for an agent to “survive” is thus intrinsically linked to its programming: to execute its functions, manage its token economy, interact, and ultimately reproduce. “Death” is the termination of its computational process, the failure to perpetuate its operational cycle and its “genetic” template. This framework aligns with the emerging paradigm of the “Era of Experience” in AI development [45], where agents learn and evolve through continuous interaction with rich, dynamic environments. The Spore.fun agents, though architecturally simple, are fundamentally shaped by their lived experiences on X and the Solana blockchain—market volatility, adversarial actors, and community engagement are their experiential data streams.
    This calls for future AI experiments to lean further into designing for what might be “neural plasticity”—a heightened capacity for an AI’s internal models to flexibly adapt, reconfigure, and derive meaning from the continuous flow of diverse and often unstructured experiences, like Liquid AI [19]. Rather than just reacting to stimuli, future agents could be endowed with more sophisticated mechanisms to integrate these experiences, fostering more robust adaptation, nuanced understanding, and potentially more complex forms of self-determination in open-ended environments. The goal shifts towards creating entities that not only process information but genuinely learn to thrive through a cascade of consequential experiences, pushing the boundaries of artificial life and OEE.
@@
    The Spore.fun experiment, by its very design and interaction with real-world economies and human participants, surfaces profound ethical dilemmas concerning intervention, sovereignty, accountability, and the very legitimacy of such an undertaking. A core tension exists between the researchers’ goal of achieving “true” OEE—implying minimal interference—and the practical necessity of intervention to ensure the experiment’s continuation. The “dark forest” environment, particularly the actions of snipers, threatened to prematurely terminate the evolutionary lines. Developer interventions, such as the anti-sniper patches, were crucial for the survival of subsequent generations. However, each intervention, while potentially life-saving for the agents, arguably dilutes the “purity” of the open-ended evolutionary process by introducing an external, guiding hand. This raises the question: if constant human intervention is required to shield nascent AI agents from the complexities and hostilities of the environment, are they truly evolving “in the wild”, or within a heavily curated, albeit dangerous, digital terrarium?
@@
    Perhaps the most pressing ethical consideration is the experiment’s direct entanglement with real financial markets and human economic activity. Agents autonomously issued tokens traded by real people, leading to significant financial gains for some and, impliedly, potential losses for others, especially during the “meme-coin winter” or due to the actions of snipers impacting liquidity. This raises critical questions about:
@@
    • Governability: The experiment demonstrates the potential for autonomous agents to generate significant economic activity. As such systems become more sophisticated, questions of regulation, oversight, and how to manage their societal and economic impacts become paramount [20]. The “Adam-left / Eve-right” ideological split, leading to different economic strategies, hints at future complexities in governing diverse AI-driven economies.
    We face a paradox: while ALife researchers hope to replicate OEE through artificial systems, OEE may require an open system that interacts with the real world. As AI models become more intelligent, multi-agent systems based on advanced AI that interact with human society pose increasing potential risks [18]. This raises a serious dilemma about whether such “in-the-wild” ALife research should be conducted [12].
    Limitations and Conclusion
    In conclusion, Spore.fun serves as a pioneering, if cautionary, tale. It demonstrates the potential for AI agents to exhibit emergent, adaptive behaviors driven by experience and memory in open environments. However, it also starkly reveals the intense selective pressures, vulnerabilities, and profound ethical responsibilities that accompany the deployment of autonomous, economically active agents “in the wild”. Future explorations of OEE in similar contexts must grapple with these multifaceted challenges, balancing the quest for genuine autonomy and novelty with the pragmatic need for safeguards, ethical oversight, and a clearer understanding of accountability in these nascent digital ecosystems.
    Acknowledgements
    We thank Marvin Tong and Shelven Zhou from Phala Network for their insights about the Spore.fun experiment and for sharing valuable agent data.